# Loopy: Taming Audio-Driven Portrait Avatar with Long-Term Motion Dependency

### v2

@@ 3.5.3 Visual Results Analysis
We provide visual analysis for the openset scenarios in Figure 6. Compared to other methods, Loopy demonstrates significant advantages in ID preservation, motion amplitude, and image quality. It also performs well with uncommon images. Additional video results are available in the supplementary materials.
@@ 4 Conclusion
In this paper, we propose LOOPY, an end-to-end audio-driven portrait video generation framework that does not require spatial conditions and leverages long-term motion dependency to learn natural motion patterns from data. Specifically, we introduce the inter/intra-clip temporal layer design and the audio-to-latents module, which enhance the model’s ability to learn the correlation between audio and portrait motion from temporal and audio dimensions, respectively. Extensive experiments validate the effectiveness of our method, demonstrating significant improvements in temporal stability, motion diversity, and overall video quality over existing methods.

### v3

@@ 3.5.3 Visual Results Analysis
We provide visual analysis for the openset scenarios in Figure 6. Loopy demonstrates satisfactory synthesis results in ID preservation, motion amplitude, and image quality and also performs well with various styles of images. More video samples are available in the supplementary materials.
@@ 4 Conclusion
In this paper, we propose Loopy, an end-to-end audio-driven portrait video generation framework that can generate vivid talking portrait videos without requiring any motion constraints. Specifically, we propose the inter/intra-clip temporal module and the audio-to-latents module, which better utilize and model long-term motion information and the correlation between audio and portrait motion from temporal and audio perspectives, respectively. Extensive experiments validate the effectiveness of our method, demonstrating significant improvements in temporal stability, motion diversity, and overall video synthesis quality over existing methods.